# Methods for differential network estimation: an empirical comparison

### v1

@@ 4 Discussion
We have investigated the performance of five different methods for differential network estimation under varying graph structure, sparsity and sample size. Overall, we find lasso penalized D-trace loss to perform the best. However, given our focus on structure learning, one clear drawback of this method is the lack of formal error control. This is an area for future research.
@@
Regarding the error control, methods for testing the equality of entries of precision matrices or partial correlations both controlled FDR at the nominal level $\alpha$, with the exception of star graph settings, where methods were much less stable and occasionally empirical FDR exceeded $\alpha$. DCA was more conservative, with empirical FDR lower then the target level in most cases. We note that DCA provides a collection of raw $p$-values that we corrected for multiplicity with the Benjamini Hochberg procedure. It is likely that a more careful consideration of the specific multiple testing issue arising here could lead to a more efficient, less conservative solution. See Figure (b) in the Supplementary materials for more details.
We found that, overall, performance was primarily related to graph structure. We would like to investigate this a bit further.
We expect that performance of methods for learning differential networks depends on various properties of underlying graphs and probability distributions. To gain further insight, various measures of these properties can be employed. We considered two measures: one that corresponds to a global distance between distributions $\mathcal{N}(0,\mathbf{\Sigma}^{(1)})$ and $\mathcal{N}(0,\mathbf{\Sigma}^{(2)})$, and one that summarises the local structure of a differential graph.
We evaluate the global distance between distributions using symmetrized Kullback-Leibler divergence $KL_{12}+KL_{21}$, where for two centered Gaussian distributions $\mathcal{N}(0,\mathbf{\Sigma}^{(i)})$, $i=1,2$, the Kullback-Leibler divergence $KL_{ij}$, $i,j=1,2$, is defined as
@@
To quantify local differences, we use the highest vertex degree of a differential graph divided by the total number of vertices $p$. Note that normalization by $p$ is not necessary in the presented simulation as all graphs have the same size, $p=200$, but we use it for possible comparison with other studies.
And finally, to summarize performance curves with a scalar, we compute the area under the curve (for power vs FDR) for $n=400$ samples and average it over all five methods. We relate this performance indicator to the symmetrized Kullback-Leibler divergence and to the highest vertex degree. The resulting plots are shown in Figure 5.
@@
The left panel of Figure 5 demonstrates that methods on average perform better for sufficiently distant distributions (e.g. the cases of random and scale-free graphs in 50-200 and 100-200 edges settings). In other words, the larger the global differences are, the easier it is to estimate them correctly. One interesting observation is that performance can vary considerably for a given distance (e.g. three points having a similar value of log2 of symmetric KL divergence $\approx$ 2.4). This, however, may be at least partially explained by the highest degree (right panel). If the highest degree is sufficiently low, performance is good despite limited global difference.
We notice that for the star differential graph, the distance between distributions is low and the normalized highest degree is high in all settings. It is well known that hub structures are harder to estimate, as most methods assume uniform distribution of edges, without any specific structure (see discussion in (Tan et al.,, 2014), for example). Therefore, more information is required for a successful estimation in these situations. Indeed, for one of the methods (testing for equality of precision matrix entries, since it is one of the fastest) we have evaluated the performance for higher samples sizes, $n_{1}=n_{2}=1000,2000$, and discovered a considerable improvement (see Figure (c) in the Supplementary materials). We assume this observation may be applicable to other methods as well. Although increased sample size may be unachievable in many situations, aiming for as many samples as possible is advisable.
@@
Methods for learning individual networks with hub structure do also exist, see e.g. the hub graphical lasso (Tan et al.,, 2014) or the hub weighted graphical lasso (McGillivray et al.,, 2020). Methods such as the perturbed node graphical lasso (Mohan et al.,, 2014) estimates two individual networks jointly taking into account potential hub differences. A possible direction of future work could be to extend such methods to the differential network situation, e.g. by using their estimands as an input for the testing-based methods described in the current paper.
@@
Finally, related to sample size, it would be desirable to obtain theoretical results on the necessary sample size $n$ in relation to the differential graph order $p$ and its maximum degree, as done, for example, for Ising models (Ravikumar et al.,, 2010).
Appendix A: Real data settings
Here, we provide the parameters used to estimate a 10-edge network in the real data example.
  • For fused graphical lasso (FGL) (Danaher et al.,, 2014), implemented in the package JGL, we set $\lambda_{1}=0.1$ (as we empirically found this value to show good performance in our simulation studies), $\lambda_{2}=0.392$ and select a differential edge if $|\hat{\Omega}_{ij}^{(1)}-\hat{\Omega}_{ij}^{(2)}|>10^{-3}$ (as the authors do in the code accompanying their manuscript);
  • For the method testing for equality of partial correlations (Liu,, 2017), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we set $\alpha=0.025$;
  • For the method testing for equality of entries of precision matrices, implemented in the package DiffNetFDR (Zhang et al.,, 2019), we set $\alpha=0.06$ (note that this result is the only one with 11 edges, as we were not able to obtain 10);
  • For the differential connectivity analysis method (Zhao and Shojaie,, 2022) implemented in the package DCA, we set $\alpha=0.02$;
  • For the direct estimation with lasso penalized D-trace loss (Yuan et al.,, 2017) (referred to as Dtrace for short in this manuscript), implemented in the package DiffGraph (Zhang et al.,, 2018), we set $\lambda=0.555$.
Appendix B: Simulation study data generation
@@
Appendix C: Simulation settings
@@
  • For the method testing for equality of partial correlations (Liu,, 2017), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we vary FDR level $\alpha$ from 0.001 to 0.995;
  • For the method testing for equality of entries of precision matrices (Xia et al.,, 2015), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we vary FDR level $\alpha$ from 0.001 to 0.995;
@@
  • For the direct estimation with lasso penalized D-trace loss (Yuan et al.,, 2017) (referred to as Dtrace for short in this manuscript), implemented in the package DiffGraph (Zhang et al.,, 2018), we vary the tuning parameter $\lambda$ from 0.1 to 0.7.
Data availability
BRCA.547.PAM50.SigClust. Subtypes.txt (cancer subtypes). The pre-processed dataset we used is a part of DiffNetFDR package (Zhang et al.,, 2019) as a TCGA.BRCA object.
Acknowledgments

### v2

@@ 5 Discussion
We have investigated the performance of six different methods for differential network estimation under varying graph structure, sparsity and sample size. We find direct estimation with lasso penalized D-trace loss to perform the best overall. However, given our focus on structure learning, one clear drawback of this method is the lack of formal error control. This is an area for future research.
@@
Regarding error control, PMat, PCor and ONDSA controlled FDR at the nominal level $\alpha$, with the exception of the star graph settings, where the methods were much less stable and occasionally empirical FDR exceeded $\alpha$ (see Figure (b) in the Supplementary materials). For ONDSA, FDR was also not controlled in the 50-200 random graph setting with 400 samples. The reason for this is not clear. Regarding the DCA, it was overall more conservative. However, in some settings empirical FDR significantly exceeded the target level for small values of $\alpha$. The inflation of FDR occured in general for the smaller sample size and was even more pronounced for the star differential graph. This finding is not surprising since error control for this method strongly relies on the coverage property, that is, on the event that the estimated neighborhood of each node covers its true neighborhood. This event has smaller probability both when $n$ is small and when the size of the neighborhood is large, as is in a star graph.
We found that overall performance was primarily related to graph structure. We would like to investigate this a bit further.
We expect that performance of methods for learning differential networks depends on various properties of underlying graphs and probability distributions. To gain further insight, various measures of these properties can be employed. We considered two measures: one that corresponds to a global distance between distributions $\mathcal{N}(0,\mathbf{\Sigma}^{(1)})$ and $\mathcal{N}(0,\mathbf{\Sigma}^{(2)})$, and one that summarizes the local structure of a differential graph.
We evaluated the global distance between distributions using symmetrized Kullback-Leibler divergence $KL_{12}+KL_{21}$, where for two centered Gaussian distributions $\mathcal{N}(0,\mathbf{\Sigma}^{(i)})$, $i=1,2$, the Kullback-Leibler divergence $KL_{ij}$, $i,j=1,2$, is defined as
@@
To quantify local differences, we used the highest vertex degree of a differential graph divided by the total number of vertices $p$. Note that normalization by $p$ is not necessary in the presented simulation as all graphs have the same size, $p=200$, but we use it for a possible comparison with other studies.
And finally, to summarize the performance curves with a scalar, we computed the area under the curve (for power vs FDR) for $n=400$ samples and averaged it over all six methods. We related this performance indicator to the symmetrized Kullback-Leibler divergence and to the highest vertex degree. The resulting plots are shown in Figure 5.
@@
The left panel of Figure 5 demonstrates that methods on average perform better for sufficiently distant distributions (e.g. the cases of random and scale-free graphs in 50-200 and 100-200 edges settings). In other words, the larger the global differences are, the easier it is to estimate them correctly.
We notice that for the star differential graph, in some settings, the distance between distributions is as high as for random and scale-free graphs, but the performance of the methods is still considerably worse, measured by mean AUC. This may be explained by the local structures, as illustrated by the normalized highest degree, that is high for the star graph in all settings. It is well known that hub structures are harder to estimate, as most methods assume a uniform distribution of edges, without any specific structure (see discussion in (Tan et al.,, 2014), for example). Therefore, more information is required for a successful estimation in these situations. Indeed, for one of the methods (PMat, since it is one of the fastest) we have evaluated the performance for higher sample sizes, $n_{1}=n_{2}=1000,2000$, and discovered a considerable improvement (see Figure (c) in the Supplementary materials). We hypothesize this observation applies to other methods as well. Although increased sample size may be unachievable in many situations, aiming for as many samples as possible is advisable.
@@
Methods for learning individual networks with hub structure do also exist, see e.g. the hub graphical lasso (Tan et al.,, 2014) or the hub weighted graphical lasso (McGillivray et al.,, 2020). Methods such as the perturbed node graphical lasso (Mohan et al.,, 2014) estimate two individual networks jointly taking into account potential hub differences. A possible direction of future work could be to extend such methods to the differential network situation, e.g. by using their estimands as an input for the testing-based methods described in the current paper.
@@
Finally, related to the sample size, it would be desirable to obtain theoretical results on the necessary sample size $n$ in relation to the differential graph order $p$ and its maximum degree, as done, for example, for Ising models (Ravikumar et al.,, 2010).
Declarations
Author contributions
AP, MT and VD conceived the research idea. AP conducted the simulations and the real data analysis, and drafted the manuscript with input from MT and VD. All authors gave final approval.
Funding
Data availability
BRCA.547.PAM50.SigClust.Subtypes.txt (cancer subtypes). The pre-processed dataset we used is a part of DiffNetFDR package Zhang et al., (2019) as a TCGA.BRCA object.
Code availability
Simulation study code is freely available at github.com/annaplaksienko/Diff_networks_review_simulation.
Appendix A Real data settings
Here, we provide the parameters used to estimate a networks with optimal parameters in the real data example.
  • For fused graphical lasso (FGL) (Danaher et al.,, 2014), implemented in the package JGL, we first fix $\lambda_{2}=0.01$ and chose $\lambda_{1}$ with BIC, obtaining $\lambda_{1}=0.55$. Then, using that value, we perform search over the grid for $\lambda_{2}$. As we fail to find BIC minimum inside the interval, we propose to chose a parameter corresponding to an ”elbow”, i.e. slight change in trend of the BIC curve. That is $\lambda_{2}=0.286$. We select a differential edge if $|\hat{\Omega}_{ij}^{(1)}-\hat{\Omega}_{ij}^{(2)}|>10^{-3}$ (as the authors do in the code accompanying their manuscript);
  • For the PCor method (Liu,, 2017), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we set $\alpha=0.05$;
  • For the PMat method (Xia et al.,, 2015), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we set $\alpha=0.05$;
  • For omics networks differential and similarity analysis (ONDSA) (Chen et al.,, 2024) implemented in the package ONDSA, we set $\alpha=0.05$;
  • For the differential connectivity analysis method (Zhao and Shojaie,, 2022), implemented in the package DCA, we set $\alpha=0.05$;
  • For the DTrace method (Yuan et al.,, 2017) implemented in the package DiffGraph (Zhang et al.,, 2018), we chose optimal parameter with BIC and obtained $\lambda=0.62$.
Appendix B Simulation study data generation
@@
Appendix C Simulation settings
@@
  • For the PCor method (Liu,, 2017), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we vary FDR level $\alpha$ from 0.001 to 0.995;
  • For the PMat method (Xia et al.,, 2015), implemented in the package DiffNetFDR (Zhang et al.,, 2019), we vary FDR level $\alpha$ from 0.001 to 0.995;
@@
  • For the DTrace method (Yuan et al.,, 2017),implemented in the package DiffGraph (Zhang et al.,, 2018), we vary the tuning parameter $\lambda$ from 0.1 to 0.7.